OpenStudy (anonymous):

Which point lies on the graph of the following piecewise function? y = x – 1 if x < 2 y = x^2 if x greater than or equal to 2 a.(1, 1) b.(2, 1) c.(–1, 1) d.(3, 9)

OpenStudy (anonymous):

d. y = x^2 since x =3 y = 3^2 y = 9 , thus the point (3,9) is on the graph

OpenStudy (anonymous):

isn't (2, 1) also on the graph?

OpenStudy (anonymous):

For point (2,1) you have to use y = x^2, you don't use y = x -1. This is because x = 2 and tells you which function to use.

OpenStudy (anonymous):

Ok, but why wouldnt you use y = x - 1, don't you use both?

OpenStudy (anonymous):

No, you only use y = x-1 when x < 2. Not when it is actually x =2. You can only use one equation per value of x.